Find the electric field produced by a uniformly polarized sphere of radius R
1
Expert's answer
2019-12-13T10:08:47-0500

The potential outside:


"V(r)=\frac{P}{3\epsilon_0}\frac{R^3}{r^2}\cos\theta=\frac{1}{4\pi \epsilon_0}\frac{\frac{4\pi PR^3}{3}}{r^2}\cos\theta"

"V(r)=\frac{1}{4\pi \epsilon_0}\frac{\bold{p\cdot \hat{r}}}{r^2}"

where

"p=\frac{4\pi PR^3}{3}"

The potential inside:


"V(r)=\frac{p}{3\epsilon_0}r\cos\theta=\frac{p}{3\epsilon_0}z"

The  electric field:


"\bold{E}=-\nabla V=-\frac{P}{3\epsilon_0}\hat{\bold{z}}"